OpenStudy (anonymous):

How much tension must a rope withstand if it is used to accelerate a 1190.0 kg car vertically upward at 0.980 m/s2

OpenStudy (michele_laino):

on you car are acting two forces: the weight force: \(W=m \cdot g= 1190 \cdot 9.81=...Newton\) and the fictitious force: \(F=m \cdot a=1190 \cdot 0.980=... Newtom\) so total force is \(W+F=...Newton\)

OpenStudy (michele_laino):